Mikaela is a STEM student who is currently having her General Chemistry 2 subject. Today, she is asked by her teacher to prepare CuSO4 solution. Help her do her task by following the procedure and answering the questions below.


1. Obtain sample of CuSO4 from the front bench.

2. Using your scale, weigh out 18g of CuSO4

3. Dissolve the 18g of CuSO4 in 100 ml of distilled water

4. Stir until you have reached a homogeneous solution with all the solute dissolved is the solvent.


Calculate the concentration of your solution (M in mol/L of CuSO4).


1
Expert's answer
2021-06-14T13:35:34-0400

Molar mass of CUSO4=159.5gm /mol

Given,

Mass of CuSO4=18gm

Volume of water =100ml=0.1L


concentration of solution in terms of molality(M)=n/v

n=number of moles

n=mass/molar mass

n=18/159.5

n=0.112mol

Now,

M=0.112/0.1

M=1.12 mol/L

Hence, the concentration in terms of molarity of CuSO4 is 1.12mol/L